OpenStudy (anonymous):

x^1over3 times x^1over3

OpenStudy (anonymous):

\[x^{1\over3} \times x^{1\over3}\]

OpenStudy (amistre64):

what are we to do?

OpenStudy (amistre64):

the expression alone is like giving the ingredients to a recipe; without giving the recipe :) eggs and butter ....

OpenStudy (anonymous):

oh sorry! haha simplify, express the product as a rational exponent! sorry I keep asking for math answers. I'm terrible at math -.-

OpenStudy (anonymous):

hahahaha nice comparison!

OpenStudy (amistre64):

when we add like bases; we add exponents. \[x^n*x^m=x^{n+m}\] \[x^{1/3}*x^{1/3}=x^{1/3+1/3}=x^{2/3}\]

OpenStudy (amistre64):

another rule of thumb is \[\Large x^{t/b}=\sqrt[b]{x^t}\] \[\Large x^{2/3}=\sqrt[3]{x^2}\]

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (amistre64):

youre welcome

OpenStudy (asnaseer):

LOL! - I had to do a double take on your question as I initially misread it as:\[x^{LOVER3}*x^{LOVER3}\]